OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10

OpenStudy (anonymous):

y = (x + 6)/3 x = (7y + 10)/2 y = ((7y + 10)/2 + 6)/3 y = (7y + 10)/6 + 2 6y=7y + 10 + 12 7y-6y = -22 y = -22 x = -72

OpenStudy (anonymous):

Solve y Not x

OpenStudy (anonymous):

Then I would say first line of Dangolbery answer : y = (x + 6)/3